OpenStudy (amina_soneviseth):

I GIVE MEDALS PLEASE HELP QUICK how to solve : -9 + (-5) x +3 + (-2)x also please show work? TY!

OpenStudy (photon336):

combine like terms first

OpenStudy (photon336):

-9-5x+3-2x group terms together. -9+3 = -6 -5x-2x = -7x -7x-6

OpenStudy (amina_soneviseth):

@Photon336 So it's -7x - 6 = ?

OpenStudy (photon336):

@amina_soneviseth what's on the other side of the problem? i noticed there was nothing on the other side of the equal sign. but yeah if we were to stop there that's what it would be.

OpenStudy (amina_soneviseth):

@Photon336 lol ik now thank you!

OpenStudy (photon336):

no problem welcome to open study